OpenStudy (anonymous):

how do can i add (2x)/(3x-5) to (x+1)/2x.. use complete sentence to explain your answer.

OpenStudy (anonymous):

Just to be clear your question is: \[\frac{2x}{3x-5}+ \frac{x+1}{2x}\] Right?

OpenStudy (anonymous):

yes @AskingTheHardQuestion

OpenStudy (anonymous):

\[\frac{2x}{3x-5}+ \frac{x+1}{2x} \implies \frac{(7x + 5) (x-1)}{2x(3x-5)}\] Step 1: Put fractions over a common denominator: 2x(3x-5).

OpenStudy (anonymous):

Step 2: Use FOIL: \[2x \times 2x(3x-5) \implies 12x-10x \implies 2x\] So far we have: \[\frac{2x}{3x-5}+\frac{x+1}{2x}\]

OpenStudy (anonymous):

Step three: \[\frac{x+1}{2 x} = \frac{4 x^2}{2 x (3 x-5)}+\frac{(x+1)(3 x-5)}{2 x (3 x-5)}\]

OpenStudy (anonymous):

Step four and five: Use FOIL, then group like terms.

OpenStudy (anonymous):

is that is?

OpenStudy (anonymous):

Well no, you could continue to FOIL this: \[\frac{(7x + 5) (x-1)}{2x(3x-5)}\] When would in the end give you: \[\frac{7 x^2-2 x-5}{6 x^2-10 x}\]

OpenStudy (anonymous):

What a long a tedious process :-P. If you wanted to expand that last part you would get: \[\frac{7 x^2}{6 x^2-10 x}-\frac{2 x}{6 x^2-10 x}-\frac{5}{6 x^2-10 x}\] Expanding is completely optional!